OpenStudy (anonymous):

Sides are added to a convex polygon so that the sum of its interior angle measures is increased by 540 degrees. How many sides are added to the polygon?

OpenStudy (anonymous):

A convex polygon has 5 sides.

OpenStudy (anonymous):

Formula for sum of interior angles.\[180(n-2)=interior sum\] Use it to find the interior sum of a polygon with 5 sides.

OpenStudy (anonymous):

\[180(n-2)=540\] Sorry find n for me. \[180n-360=540\]